OpenStudy (anonymous):

Let f(x) = 6x^{3}+7. Find the open intervals on which f is increasing (decreasing). Then determine the x-coordinates of all relative maxima (minima). 1. f is increasing on the intervals 2. f is decreasing on the intervals 3. The relative maxima of f occur at x = 4. The relative minima of f occur at x =

OpenStudy (freckles):

so this is a cal question?

OpenStudy (anonymous):

yes it is. calc 1

OpenStudy (freckles):

ok first step to get any of this is to differentiate f

OpenStudy (freckles):

\[f(x)=6x^{3}+7\] so \[f'(x)\]

OpenStudy (freckles):

there is suppose to be =? after the f'(x) there

OpenStudy (anonymous):

f'(x) = 18x^2

OpenStudy (freckles):

now if f'>0 then f is increasing if f'<0 then f is decreasing

OpenStudy (freckles):

when f'=0 we have f is doing neither and f'=0 gives us our critical numbers

OpenStudy (freckles):

now what do you know about the factor x^2? Is x^2 ever negative?

OpenStudy (anonymous):

it is never negative

OpenStudy (freckles):

\[\text{ since } 18x^2 \ge 0 \text{ for all x} \\ \text{ then this tells us } \text{ f is never descreasing }\]

OpenStudy (freckles):

since f' is never negative

OpenStudy (freckles):

anyways 18x^2=0 when x=?

OpenStudy (anonymous):

x = 0

OpenStudy (freckles):

so that means everywhere but at x=0 f is increasing since f'=18x^2>=0 for all x

OpenStudy (freckles):

also a function that is never decreasing doesn't have any relative min or max just like a function that is never increasing doesn't have any relative min or max for there to be a min or max at a critical number the function must switch from decreasing to increasing there or increasing to decreasing

OpenStudy (freckles):

that is for a relative min/ max anyways

OpenStudy (anonymous):

okay so there is no relative min or max, nor is the function have any decreasing intervals. So it only has an increasing interval at (-inf,inf). Is this correct?

OpenStudy (freckles):

well i would say isn't increasing at x=0

OpenStudy (freckles):

because f'=0 at x=0

OpenStudy (freckles):

but it is increasing everywhere else

OpenStudy (freckles):

(-inf,0) U (0,inf)

OpenStudy (anonymous):

Okay, thanks for the help.

OpenStudy (freckles):

np